# CoT-Space: A Theoretical Framework for Internal Slow-Thinking via Reinforcement Learning

### v1

@@ 3.1 Reasoning as Optimization: A Perspective of Noise
Building on this, we can model the optimization of LLM reasoning as a form of reasoning loss descent and perform a theoretical analysis. As previously established, when viewed from a reasoning-level perspective, the reasoning space can be considered approximately continuous, i.e., $s_{o}\in\mathcal{S}$ where $\mathcal{S}$ is a continuous space. The state is then optimized continuously based on the policy and reasoning loss. This perspective enables us to analyze the reasoning process through a loss-reduction lens.
@@
Equations (1) and (2) formulate the optimization process of CoT reasoning from the gradient update and SDE perspective, respectively. By combining the two equations related to noise, we can thus obtain the relationship between noise scale and CoT length, as shown in the following theorem.
Theorem 3.1.
The proof is provided in Appendix A.2. Theorem 3.1 clarifies the relationship between the noise $g$ and the CoT length $L$ in the same the reasoning process. According to the sharpness-aware minimization theorems, given a reasoning task, there exists an optimal noise scale $g_{\text{opt}}$ in the reasoning process, which corresponds to an optimal CoT length $L_{\text{opt}}\propto\frac{1}{g_{\text{opt}}}$. In essence, this theoretical analysis provides a fundamental basis for why an optimal CoT length is crucial for effective reasoning, highlighting its role in introducing the right amount of noise to achieve a flat, high-quality, and generalizable solution.
3.2 Reasoning and Generalization: A Perspective of Risk
The previous section considered reasoning from the perspective of noise and analyzed it through an optimization lens. Here, we analyze the generalization risk of a policy to provide further theoretical insights, focusing on the upper bounds of reasoning risks.
In the context of CoT-Space, risks can be defined similarly to classical ML. The total expected error of a policy $\pi$ on the true data distribution $\mathcal{D}$, denoted by $R(\pi)=\mathbb{E}_{q\sim\mathcal{D}}[C(s_{q}^{\pi})]$, can be decomposed into two distinct components:
$R(\pi)=\underbrace{\hat{R}_{n}(\pi)}_{\text{Empirical Loss (Bias)}}+\underbrace{\left(R(\pi)-\hat{R}_{n}(\pi)\right)}_{\text{Generalization Error (Variance)}}.$ (3)
Here, $\hat{R}_{n}(\pi)=\frac{1}{n}\sum_{i=1}^{n}C(s_{q_{i}}^{\pi})$ is the empirical loss on the training set. In the following parts, we will analyze how each component behaves as a function of the CoT complexity, which we proxy by the expected CoT length $L$.
3.2.1 An Upper Bound for Generalization Error: The Overfitting Risk
We define the reasoning generalization error as the difference between the true expected loss and the empirical loss. This error quantifies a policy’s ability to generalize its reasoning to unseen queries, a concept parallel to the generalization error in traditional machine learning theory.
Definition 3.2.
(Reasoning generalization error.) Given $n$ queries $\{q_{i}\}_{i=1}^{n}$ sampled i.i.d. from a distribution $\mathcal{D}$, and a policy $\pi$ learned from $\{q_{i}\}_{i=1}^{n}$, let’s denote $s_{q}^{\pi}=\pi(q)$ as the end reasoning state output by $\pi$ on query $q$. The reasoning generalization error of $\pi$ on $\mathcal{Q}$ is defined as:
$R(\pi)-\hat{R}_{n}(\pi)=\mathbb{E}_{q\sim\mathcal{D}}[C(s_{q}^{\pi})]-\frac{1}{n}\sum_{i=1}^{n}C(s_{q_{i}}^{\pi}).$
To derive an upper bound for this error in Definition 3.2, we can simply adapt the PAC-Bayes framework, which provides a powerful link between generalization, policy complexity, and data volume. For our analysis within CoT-Space, we establish the following components:
By applying the PAC-Bayes theorem [18], we obtain a bound on the expected true loss which, when rearranged, yields an upper bound on the expected generalization error.
Theorem 3.3.
$\mathbb{E}_{\pi\sim Q}[R(\pi)-\hat{R}_{n}(\pi)]\leq\sqrt{\frac{\operatorname{KL}(Q||P)+\ln\frac{2n}{\delta}}{2n}}$
A detailed analysis is provided in Appendix A.3. Theorem 3.3 provides a powerful, high-probability upper bound on the expected reasoning generalization error for policies sampled from the learned posterior $Q$. While this PAC-Bayesian bound is a valid, generic result, we can derive a more specific bound by using information theory, a perspective that connects the generalization error to the mutual information between the learned policy and the training data. This approach leads to the following information-theoretic upper bound.
Theorem 3.4.
A brief proof is provided in Appendix A.4. Theorem 3.4 explicitly connects the generalization error to the observable, structural properties of the reasoning process. It suggests that overfitting, or “overthinking”, can occur in two distinct ways: through excessive reasoning depth (high $\mathbb{E}[L]$) or through excessive step verbosity (high $\mathbb{E}[|\xi|]$), both resulting in overlong response length. This provides a direct theoretical basis for regularizing the CoT generation process in RL to improve reasoning generalization.
3.2.2 A Lower Bound for Empirical Loss: The Underfitting Risk
Definition 3.5.
Definition 3.5 can be likened to the minimum number of steps required to solve a reasoning problem. For instance, a simple two-digit addition problem might require only one or two mental steps, making its $L^{*}$ very low. In contrast, proving a complex theorem, such as the Pythagorean theorem, requires a sequence of many interconnected logical steps, resulting in a high $L^{*}$. A policy that attempts to solve the theorem with only a few steps, regardless of how smart it is, will inevitably fail.
Assumption 3.6.
Assumption 3.6 claims that a policy’s failure due to insufficient depth can be observed in practice. Consider a complex physics problem requiring multiple steps: identifying the forces, setting up equations, and solving for variables. A policy that only completes the first step and then stops will not be able to reach the correct final answer. In this case, the loss function would assign a high penalty (our $C_{fail}$) because the generated CoT, while possibly correct for the initial step, is fundamentally incomplete for solving the problem. The policy, therefore, cannot even fit the training data for problems that require deeper reasoning.
With this framework established by Definition 3.5 and Assumption 3.6, we can derive a lower bound for the empirical reasoning loss $\hat{R}_{n}(\pi)$.
Theorem 3.7.
$\hat{R}_{n}(\pi)\geq P_{\pi}(L<L^{*})\cdot C_{fail},$
where $P_{\pi}(L<L^{*})=n_{fail}/n$ is the empirical failure rate of policy $\pi$ on the training set.
A detailed proof is provided in Appendix A.5. Theorem 3.7 provides a formal basis for the “underfitting” portion of our trade-off curve. It demonstrates that the empirical loss is directly constrained by the probability that the policy generates a CoT that is too short for the problem’s intrinsic complexity. A policy that tends to produce short CoTs (i.e., has a small $\mathbb{E}[L]$) will have a high failure rate $P_{\pi}(L<L^{*})$ on any dataset containing non-trivial problems. Consequently, its empirical loss $\hat{R}_{n}(\pi)$ is guaranteed to be high. This formalizes why a policy must learn to increase its CoT length $L$ to achieve a low empirical loss, thus completing our explanation for the tradeoff in our total error analysis and the existence of an optimal $L_{\text{opt}}$.
3.2.3 Tradeoff between Overfitting and Underfitting Risk
In Theorem 3.4, we provide an upper bound for the generalization error in the reasoning process with relation to CoT length $L$. While in Theorem 3.7, we provide a lower bound for the empirical loss in the reasoning process with relation to CoT length $L$.
Since the total expected error $R(\pi)$ is the sum of a decreasing function (empirical loss) and an increasing function (generalization error), this classic Bias-Variance Trade-off inevitably leads to a U-shaped curve for the total error as a function of the CoT length $L$. For small $L$, the total error is high due to high bias (underfitting). The policy can’t solve even the training problems. For large $L$, the total error is high due to high variance (overfitting). The policy has too much capacity and learns non-generalizable details.
Therefore, an optimal length, $L_{\text{opt}}$, must exist that minimizes the total error. This $L_{\text{opt}}$ is long enough to avoid underfitting but short enough to prevent significant overfitting. Our analysis reveals that there is no universal best CoT length, instead, the existence of an optimal length is a natural consequence of the fundamental trade-off between fitting the training data and generalizing to new data. From this perspective, any RL algorithm that can explore effectively can eventually converge to this optimal strategy.
3.3 Takeaways
We have analyzed with the CoT-Space from two perspectives: that of noise (Subsection 3.1) and risk (Subsection 3.2). Based on the theoretical insights from Theorems 3.1, 3.4, and 3.7, we can draw the following key conclusions about the factors that govern the convergence of the optimal CoT length $L_{\text{opt}}$ in LLM reasoning via RL.
Remark 1. The intrinsic difficulty of a reasoning task is a primary determinant of the optimal CoT length. According to Theorem 3.7, a more challenging task necessitates a greater $L^{*}$ to achieve a low empirical loss and avoid underfitting. Consequently, the optimal CoT length, $L_{\text{opt}}$, must be sufficiently large to accommodate the task’s complexity, otherwise the policy’s failure rate will remain high. This insight formalizes the intuitive idea that complex problems require more thinking steps to be solved correctly, and any effective policy must learn to generate a CoT of a length commensurate with the problem’s difficulty.
Remark 2. The capacity of the policy model directly influences the optimal CoT length by affecting the risk of overfitting. A model with higher capacity (e.g., more parameters) can explore a more expansive reasoning state space. This expanded exploration enables it to access states with higher reasoning losses that are inaccessible to weaker models, leading to a larger effective $C_{max}$. The generalization bound in Theorem 3.4 is an increasing function of both $C_{max}$ and the CoT length $L$. Consequently, to mitigate the heightened overfitting risk posed by its own high capacity and the resulting larger effective $C_{max}$, a more powerful model must learn to regularize its output. It achieves this by converging to a smaller optimal CoT length ($L_{\text{opt}}$), thereby maintaining a low generalization error.
Remark 3. While the choice of RL algorithm is crucial for effective training, our analysis suggests it has a limited direct influence on the final, converged CoT length. The optimal CoT length ($L_{\text{opt}}$) is primarily a function of the policy’s capacity and the inherent complexity of the training data. The RL algorithm’s main role is to find a policy that converges to this optimal length. As long as the algorithm is effective in its optimization, the final $L_{\text{opt}}$ is a property of the learned policy and the data distribution, rather than a byproduct of the specific RL approach used.
Remark 4. In a given problem space, the scale of noise in the optimization process is inversely correlated with the converged CoT length. According to Theorem 3.1, when a policy is exploring the same task space, a higher level of noise will lead to a smaller converged optimal CoT length ($L_{\text{opt}}$). This implies that a noisier optimization environment tends to favor simpler, shorter reasoning paths, likely because longer CoTs are more susceptible to the disruptive effects of noise. This finding underscores the importance of a stable training environment for achieving complex and in-depth reasoning.
To substantiate these theoretical derivations and demonstrate their practical relevance, we will now proceed to a series of empirical investigations in the following section. These experiments are designed to systematically test the four central remarks, providing robust empirical evidence for our proposed framework.
4 Experiments
This section presents a series of experiments designed to empirically validate the theoretical insights established in the preceding sections. After detailing the experimental setup in Subsection 4.1, we systematically investigate the relationship between the optimal CoT length ($L_{\text{opt}}$) and key factors, including task difficulty, model capacity, the choice of RL algorithm, and optimization noise scale in Subsections 4.2 through 4.5, respectively.
4.1 Experimental Settings
Dataset The experiments are conducted on the MATH [19] and GSM8K [20] benchmarks. The MATH training set is manually subdivided into five difficulty levels (Level 1–5) according to the original “level” field, where Level 5 is substantially more challenging than Level 1.
  • Qwen2.5 series [21]: Qwen2.5-7B, Qwen2.5-3B, Qwen2.5-1.5B, and Qwen2.5-0.6B.
  • Qwen3 series [22]: Qwen3-8B-Base, Qwen3-4B-Base, Qwen3-1.7B-Base, and Qwen3-0.6B-Base.
  • Llama3 series [23]: Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, and Llama-3.2-1B-Instruct.
Algorithm and Hyperparameters The evaluation covers the current mainstream reinforcement-learning algorithms for mathematical reasoning: GRPO [13], PPO [24], DAPO [15], Reinforce++ [25], and RLOO [26]. The reward function is rule-based [9]. For PPO, the critic model is initialized from Qwen2.5-0.6B. The learning rate for critic training is $1\times 10^{-5}$, the k3 KL-regularization coefficient is 0.001, and the GAE parameters $\lambda$ and $\gamma$ are both set to $1$. For DAPO, no KL regularization is applied; the high and low clip ratios are $0.28$ and $0.20$, respectively, and $L_{cache}$ is $1,024$. Across all algorithms, each training sample is sampled eight times ($K=8$). The actor is optimized with AdamW [27] at a learning rate of 1e-6, the training batch size is $128$, and training proceeds in a purely on-policy manner for up to 50 steps. All experiments are run on 8 $\times$ 96 GB NVIDIA H20 GPUs.
Figure 4: Empirical Validation of the Factors Governing Optimal CoT Length ($L_{\text{opt}}$). (a) The relationship between task difficulty and $L_{\text{opt}}$, demonstrating that more complex tasks require longer reasoning process. (b) The inverse relationship between model capacity and $L_{\text{opt}}$, where more powerful models converge to shorter CoTs to mitigate overfitting. (c) The convergence of mean response length under different RL algorithms, showing that the final $L_{\text{opt}}$ is largely algorithm-agnostic. (d) The inverse correlation between optimization noise scale (proxied by KL-Loss) and $L_{\text{opt}}$, indicates that noisier training environments favor more concise reasoning.
4.2 Task Difficulty and Converged Length
To empirically validate Remark 1, we investigate the relationship between task difficulty and the converged optimal CoT length ($L_{\text{opt}}$). In this experiment, we utilize the converged mean response length of the policy (Qwen2.5-7B-Base) as a direct empirical measure for $L_{\text{opt}}$. The intrinsic difficulty of the task is stratified by using datasets of increasing complexity, namely GSM8K and the five manually curated levels of the MATH dataset (Level 1 to Level 5). As illustrated in Figure 4(a), there is a clear and monotonic trend: as the complexity of the task increases, the corresponding $L_{\text{opt}}$ also increases significantly. This result provides strong evidence for our theoretical claim. According to Theorem 3.7, more challenging tasks possess a greater required reasoning depth ($L^{*}$). Consequently, the policy must generate a longer CoT to avoid the high empirical loss associated with underfitting. The experimental outcome confirms that the RL process successfully identifies and converges to a longer $L_{\text{opt}}$ to accommodate the inherent complexity of more difficult problems. We also provide validations on more base models in Appendix B.
4.3 Model Capacity and Converged Length
This subsection examines the influence of model capacity on the optimal CoT length, thereby verifying Remark 2. We train the models by GRPO, then use the parameter count of different models as a proxy for their capacity and the final critic score as a measure of their reasoning performance. Figure 4(b) plots the converged CoT length ($L_{\text{opt}}$) against this critic score for various models trained on MATH. The results reveal a distinct negative correlation: within each model family (e.g., Llama 3, Qwen 2.5, and Qwen 3), larger and higher-performing models consistently converge to a shorter $L_{\text{opt}}$. This empirical finding aligns precisely with our theoretical analysis. As stated in Remark 2 and derived from the generalization bound in Theorem 3.4, models with higher capacity are inherently more susceptible to overfitting. The risk of overfitting is amplified by longer and more complex CoTs. Therefore, to minimize the total reasoning error, a more powerful model learns to regularize its reasoning process by adopting a more concise, shorter CoT, effectively balancing the trade-off between empirical performance and generalization.
4.4 RL Algorithm and Converged Length
In this experiment, we assess the impact of different RL algorithms on the final converged CoT length to test the validity of Remark 3. We track the mean response length over the course of training for five different RL algorithms: REINFORCE++, RLOO, PPO, DAPO, and GRPO. The resulting dynamics are displayed in Figure 4(c). While the convergence trajectories and learning speeds differ markedly among the algorithms—for example, GRPO gradually increases the response length, whereas RLOO maintains a high length from the outset—all tested algorithms ultimately steer the policy toward a similar range of converged mean response lengths for a given task and model. This outcome supports our insight in Remark 3 that the choice of RL algorithm is primarily a facilitator of the optimization process rather than a determinant of the optimal solution’s properties. The final converged length, $L_{\text{opt}}$, appears to be an intrinsic characteristic dictated by the interplay between model capacity and task difficulty, not an artifact of the specific optimization algorithm used to find it.
4.5 Noise Scale and Converged Length
Finally, we validate Remark 4 by analyzing the relationship between the converged CoT length and the noise scale of the optimization process. For this analysis, we employ the Kullback-Leibler (KL) divergence loss, or “KL-Loss”, as a quantifiable proxy for optimization noise. A higher KL-Loss indicates a greater deviation from the reference policy, representing a less stable and thus “noisier” update. Figure 4(d) illustrates the relationship between the final KL-Loss and the converged length ($L_{\text{opt}}$) across various models and task difficulties. The data shows a clear inverse correlation: for a given task, models that exhibit a higher KL-Loss consistently converge to a shorter $L_{\text{opt}}$. This empirical result substantiates the relationship derived in Theorem 3.1. A noisier optimization environment penalizes longer, more intricate reasoning paths, as they are more susceptible to disruption from unstable policy updates. Consequently, the RL process naturally favors shorter and more robust CoTs to ensure stable convergence, confirming that a higher noise scale leads to a smaller optimal CoT length.
5 Related Work
5.1 Reasoning with LLMs
Recent advancements in Large Language Models (LLMs) have significantly enhanced their capacity for complex reasoning. Building upon the foundational concept of LLMs as few-shot learners [28], a pivotal development has been the use of prompting techniques to elicit explicit reasoning processes. Chain-of-Thought (CoT) prompting [1], for instance, guides models to generate intermediate steps, a method whose robustness was later improved by Self-Consistency through the aggregation of multiple reasoning paths [8]. Theoretical understanding has advanced in parallel, with information-theoretic frameworks offering methods to quantify and mitigate reasoning errors, thus bridging empirical and theoretical work [29]. Concurrently, explorations into the representation of reasoning have yielded novel formats, from structured code in Program-of-Thought [30] to internal thought tokens generated within the model’s latent space [31, 32]. However, these explorations have been largely empirical, often lacking a strong theoretical foundation.
5.2 RL for LLM Reasoning
Test-time scaling has emerged as an effective technique for enhancing the reasoning capabilities of Large Language Models (LLMs) [33], with its underlying inference scaling laws having been empirically validated [34]. Reinforcement Learning (RL) represents a prominent methodology for implementing these scaling techniques [2]. The application of RL to LLMs has spurred two parallel lines of inquiry. The first focuses on the development of novel frameworks for post-training, such as VC-PPO [14], DAPO [15], and VAPO [16]. The second line of inquiry seeks to elucidate the fundamental impact of RL on LLM behavior, where researchers have explored diverse questions. For instance, Setlur et al. [35] demonstrated the conditional advantages of RL over Supervised Fine-Tuning (SFT). Yue et al. [36] empirically investigated whether RL can expand the knowledge boundaries of LLMs, while Shao et al. [37] examined the influence of reward modeling. Concurrently, Zhao et al. [38] posited that RL primarily serves to amplify behaviors acquired during pre-training. Although these studies provide valuable insights from different perspectives, the absence of a common theoretical formulation can lead to divergent conclusions.
5.3 Optimization and Generalization
Optimization in machine learning refers to the process of finding the minima of a loss function during model training. A prominent framework in this domain is Sharpness-Aware Minimization (SAM), which operates on the principle that flatter minima correlate with better generalization [39]. This phenomenon was later attributed to the inherent noise within the optimization process [40]. Building on this insight, subsequent works have proposed various techniques to enhance optimization, such as batch augmentation [41], extrapolation methods [42], and the introduction of a surrogate gap [43]. Generalization analysis evaluates a model’s ability to perform effectively on unseen data. Information theory provides a powerful theoretical foundation for this analysis. Key concepts include entropy, $H(X)$, which measures the information content of a random variable $X$, and mutual information, I(X;Y)=H(X)H(X|Y)I(X;Y)=H(X)-H(X|Y), which quantifies the information shared between two variables. Information-theoretic approaches have found wide-ranging applications in machine learning, from establishing bounds on the generalization capacity of deep learning models [44, 45] to enhancing model interpretability [46, 47, 48]. More recently, these tools have been employed to analyze synthetic data generation [49] and to quantify reasoning errors in LLMs [29].
6 Conclusion
In this paper, we addressed the critical misalignment between the token-level formulation of traditional reinforcement learning theories and the reasoning-level execution of Chain-of-Thought in Large Language Models. We introduced CoT-Space, a new theoretical framework that shifts the perspective from a discrete Markov Decision Process to a continuous optimization problem in a semantic reasoning space. This novel viewpoint models the generation of a reasoning chain as a trajectory towards a solution minimum in a reasoning loss landscape, providing a more intuitive and powerful lens for theoretical analysis.
Our analysis, conducted within the CoT-Space, validates the effectiveness of this framework. From an optimization perspective, we established an analogy between CoT length and the learning rate in classical machine learning, demonstrating that an optimal reasoning length corresponds to an optimal noise scale that promotes convergence to flatter, more generalizable solutions. Furthermore, from a risk analysis perspective, we decomposed the total reasoning error into empirical loss and generalization error. We proved that there exists a fundamental trade-off: shorter CoTs risk underfitting due to insufficient reasoning depth, while longer CoTs risk overfitting by increasing model complexity. This trade-off theoretically grounding the existence of an optimal CoT length $L_{opt}$, that minimizes total error.
Looking forward, the CoT-Space framework opens up several promising avenues for future research. First, it can be extended to analyze the impact of different RL algorithms and reward structures on the reasoning process, potentially leading to more principled algorithm design. Second, the insights from our risk analysis could inspire new regularization techniques aimed at explicitly controlling CoT complexity to prevent overfitting and improve generalization. Finally, this framework could be applied to other forms of complex reasoning beyond CoT, such as tree-based or graph-based thought processes, thereby advancing our theoretical understanding and practical ability to build more capable and reliable AI reasoning systems.
Appendix A Proofs
A.1 Proof of Theorem 2.6
Proof.
First, consider a single, fixed reasoning-level path $(\xi_{1},...,\xi_{L})$, where each step $\xi_{l}$ is realized by an average of $k$ tokens. By Assumption 2.5, the number of token-level realizations for this single path, $N_{\text{realize}}$, is bounded below:
$N_{\text{realize}}\geq\prod_{l=1}^{L}c^{k}=(c^{k})^{L}=c^{kL}.$ (4)
Next, we consider the total number of distinct, valid reasoning states, $|\mathcal{S}_{\text{reasoning}}^{(K)}|$. For complex problems, the number of distinct valid reasoning paths itself grows with $K$. As each additional reasoning step can potentially branch from multiple previous states, the number of reachable states grows super-linearly, often exponentially, with $K$. This implies $|\mathcal{S}_{\text{reasoning}}^{(K)}|\propto\alpha^{K}$ for some base $\alpha>1$.
From Definition 2.4, the reasoning state density $\rho(K)$ is thus:
$\rho(L)=\frac{|\mathcal{S}_{\text{reasoning}}^{(K)}|}{|\mathcal{V}_{\text{semantic}}|}\propto\frac{\alpha^{K}}{|\mathcal{V}_{\text{semantic}}|}.$ (5)
Assuming the semantic volume $\mathcal{V}_{\text{semantic}}$ for a given problem domain is fixed, the state density $\rho(K)$ grows exponentially with $K$.
Finally, we connect high density to the continuum property. In a $D$-dimensional space, the expected distance $\mathbb{E}[d(s_{i},s_{j})]$ between adjacent points $s_{i},s_{j}$ is inversely related to the density, approximately scaling as $\mathbb{E}[d(s_{i},s_{j})]\propto\rho(K)^{-1/D}$. Since $\rho(K)$ grows exponentially, it follows that:
$\lim_{K\to\infty}\mathbb{E}[d(s_{i},s_{j})]=0.$ (6)
When the expected distance between states approaches zero, it implies that for any state and any direction, another valid state can be found within an infinitesimally small neighborhood. This is the defining characteristic of a continuous space or manifold. Therefore, as the reasoning token amount $K$ increases, the discrete CoT-Space increasingly resembles a continuum, which validates the application of continuous optimization tools for its analysis. ∎
A.2 Proof of Theorem 3.1
Proof.
@@ Appendix A Proofs
This result shows that the noise scale $g$ in our CoT-Space SDE model is inversely proportional to the CoT length $L$. ∎
A.3 Proof and Analysis of Theorem 3.3
@@
Lemma A.1.
(PAC-Bayes generalization upper bound [18].) For any loss function bounded in $[0,1]$, for any prior $P$, with probability at least $1-\delta$ over the draw of the training set $\{q_{i}\}_{i=1}^{n}$, the following holds for all posterior distributions $Q$:
@@
Let’s adapt Lemma A.1 to our framework. We first need to normalize our reasoning loss $C(\cdot)$ to be in $[0,1]$. Assuming there is a maximum possible loss $C_{max}$, we can define a normalized loss $C^{\prime}(\cdot)=C(\cdot)/C_{max}$. The bound then applies to the normalized losses $R^{\prime}(\pi)$ and $\hat{R}^{\prime}_{n}(\pi)$. For simplicity of notation, let’s assume $C(\cdot)$ is already normalized in $[0,1]$. ∎
Information-Theoretic Interpretation
@@
  1. 1.
    Complexity Term ($\operatorname{KL}(Q||P)$): The KL divergence term measures the information gain in moving from our prior belief $P$ to our posterior belief $Q$ after seeing the data. It quantifies the complexity of the learned policy distribution.
    • If $Q$ is very different from $P$ (i.e., $\operatorname{KL}(Q||P)$ is large), it means we had to learn a great deal from the training data to find a good policy. This is analogous to overfitting or memorizing the training set. The bound becomes looser (larger), penalizing this complexity and suggesting a higher risk of poor generalization.
    • If $Q$ remains close to $P$ ($\operatorname{KL}(Q||P)$ is small), it implies the learned policy did not stray far from our initial bias towards simpler policies. The bound is tighter, suggesting better generalization.
  2. 2.
@@
In conclusion, this PAC-Bayes bound formalizes the trade-off in learning a reasoning policy. To achieve good generalization, a policy $\pi$ must not only achieve a low empirical reasoning loss $\hat{R}_{n}(\pi)$ on the queries it has seen, but it must also remain "simple" in an information-theoretic sense, by not deviating too much from a reasonable prior. This framework provides a principled way to analyze and potentially control for phenomena like "overthinking" by viewing it as a form of overfitting, which would correspond to a large $\operatorname{KL}(Q||P)$ term.
A.4 Proof of Theorem 3.4
@@
Lemma A.2.
(Information-theoretical generalization upper bound [45].) With CoT-Space defined above, if the policy $\pi$ is trained on an i.i.d. drawn training set $S$ with size $n$, the following upper bound for the reasoning generalization error holds:
@@
Lemma A.2 transforms the problem of bounding generalization error into the problem of bounding the mutual information $I(\pi;S)$. To make this bound meaningful for our framework, we must connect $I(\pi;S)$ to the structural properties of the generated Chain-of-Thought.
@@
Lemma A.3 formalizes the intuition that the information a policy learns from data cannot exceed the information content it is capable of expressing in its output. Together with Lemma A.2, this finishes the proof. ∎
A.5 Proof of Theorem 3.7
@@
Appendix B More Experiments

### v2

@@ Appendix B Proofs
As previously established, when viewed from a reasoning-level perspective, the reasoning space can be considered approximately continuous, i.e., $s_{o}\in\mathcal{S}$ where $\mathcal{S}$ is a continuous space. The state is then optimized continuously based on the policy and reasoning loss. This perspective enables us to analyze the reasoning process through a loss-reduction lens.
@@
Equations (29) and (30) formulate the optimization process of CoT reasoning from the gradient update and SDE perspective, respectively. By combining the two equations related to noise, we can thus obtain the relationship between noise scale and CoT length, as shown in Theorem 3.1. The detailed derivation of this result is provided below.
@@
This result shows that the noise scale $g$ in our CoT-Space SDE model is inversely proportional to the CoT length $L$.
This finishes the proof. ∎
B.5 Proof and Analysis of Theorem 3.2
To derive an upper bound for the reasoning generalization error, we can first simply adapt the PAC-Bayes framework, which provides a powerful link between generalization, policy complexity, and data volume. For our analysis within CoT-Space, we establish the following components:
By applying the PAC-Bayes theorem (McAllester, 1998), we obtain a bound on the expected true loss which, when rearranged, yields an upper bound on the expected generalization error.
Theorem B.8.
$\mathbb{E}_{\pi\sim Q}[R(\pi)-\hat{R}_{n}(\pi)]\leq\sqrt{\frac{\operatorname{KL}(Q||P)+\ln\frac{2n}{\delta}}{2n}},$
@@
Lemma B.9.
(PAC-Bayes generalization upper bound (McAllester, 1998).) For any loss function bounded in $[0,1]$, for any prior $P$, with probability at least $1-\delta$ over the draw of the training set $\{q_{i}\}_{i=1}^{n}$, the following holds for all posterior distributions $Q$:
@@
Let’s adapt Lemma B.9 to our framework. We first need to normalize our reasoning loss $C(\cdot)$ to be in $[0,1]$. Assuming there is a maximum possible loss $C_{max}$, we can define a normalized loss $C^{\prime}(\cdot)=C(\cdot)/C_{max}$. The bound then applies to the normalized losses $R^{\prime}(\pi)$ and $\hat{R}^{\prime}_{n}(\pi)$. For simplicity of notation, let’s assume $C(\cdot)$ is already normalized in $[0,1]$.
This finishes the proof. ∎
@@
Complexity Term ($\operatorname{KL}(Q||P)$): The KL divergence term measures the information gain in moving from our prior belief $P$ to our posterior belief $Q$ after seeing the data. It quantifies the complexity of the learned policy distribution. If $Q$ is very different from $P$ (i.e., $\operatorname{KL}(Q||P)$ is large), it means we had to learn a great deal from the training data to find a good policy. This is analogous to overfitting or memorizing the training set. The bound becomes looser (larger), penalizing this complexity and suggesting a higher risk of poor generalization. If $Q$ remains close to $P$ ($\operatorname{KL}(Q||P)$ is small), it implies the learned policy did not stray far from our initial bias towards simpler policies. The bound is tighter, suggesting better generalization.
@@
In conclusion, this PAC-Bayes bound formalizes the trade-off in learning a reasoning policy. To achieve good generalization, a policy $\pi$ must not only achieve a low empirical reasoning loss $\hat{R}_{n}(\pi)$ on the queries it has seen, but it must also remain simple in an information-theoretic sense, by not deviating too much from a reasonable prior. This framework provides a principled way to analyze and potentially control for phenomena like "overthinking" by viewing it as a form of overfitting, which would correspond to a large $\operatorname{KL}(Q||P)$ term.
Theorem B.8 provides a powerful, high-probability upper bound on the expected reasoning generalization error for policies sampled from the learned posterior $Q$. While this PAC-Bayesian bound is a valid, generic result, we can derive a more specific bound by using information theory, a perspective that connects the generalization error to the mutual information between the learned policy and the training data. This approach leads to the information-theoretic upper bound in Theorem 3.2.
We continue to provide the proof of Theorem 3.2 here.
@@
Lemma B.10.
(Information-theoretical generalization upper bound (Xu & Raginsky, 2017).) With CoT-Space defined above, if the policy $\pi$ is trained on an i.i.d. drawn training set $S$ with size $n$, the following upper bound for the reasoning generalization error holds:
@@
Lemma B.10 transforms the problem of bounding generalization error into the problem of bounding the mutual information $I(\pi;S)$. To make this bound meaningful for our framework, we must connect $I(\pi;S)$ to the structural properties of the generated Chain-of-Thought.
@@
Lemma B.11 formalizes the intuition that the information a policy learns from data cannot exceed the information content it is capable of expressing in its output.
Together with Lemma B.10, this finishes the proof. ∎
B.6 Proof of Theorem 3.3
@@
Definition B.12.
Definition B.12 can be likened to the minimum number of steps required to solve a reasoning problem. For instance, a simple two-digit addition problem might require only one or two mental steps, making its $L^{*}$ very low. In contrast, proving a complex theorem, such as the Pythagorean theorem, requires a sequence of many interconnected logical steps, resulting in a high $L^{*}$. A policy that attempts to solve the theorem with only a few steps, regardless of how smart it is, will inevitably fail.
Assumption B.13.
Assumption B.13 claims that a policy’s failure due to insufficient depth can be observed in practice. Consider a complex physics problem requiring multiple steps: identifying the forces, setting up equations, and solving for variables. A policy that only completes the first step and then stops will not be able to reach the correct final answer. In this case, the loss function would assign a high penalty (our $C_{fail}$) because the generated CoT, while possibly correct for the initial step, is fundamentally incomplete for solving the problem. The policy, therefore, cannot even fit the training data for problems that require deeper reasoning.
Building upon Definition B.12 and Assumption B.13, we can then derive a lower bound of the empirical reasoning loss $\hat{R}_{n}(\pi)$.
@@
Appendix C Experimental Settings
Dataset The experiments are conducted on the MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021) benchmarks. The MATH training set is manually subdivided into five difficulty levels (Level 1–5) according to the original “level” field, where Level 5 is substantially more challenging than Level 1.
  • Qwen2.5 series (Qwen, 2025a): Qwen2.5-7B, Qwen2.5-3B, Qwen2.5-1.5B, and Qwen2.5-0.6B.
  • Qwen3 series (Qwen, 2025b): Qwen3-8B-Base, Qwen3-4B-Base, Qwen3-1.7B-Base, and Qwen3-0.6B-Base.
  • Llama3 series (Meta, 2024): Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, and Llama-3.2-1B-Instruct.
Algorithm and Hyperparameters The evaluation covers the current mainstream reinforcement-learning algorithms for mathematical reasoning: GRPO (Shao et al., 2024), PPO (Schulman et al., 2017), DAPO (Yu et al., 2025), Reinforce++ (Hu et al., 2025), and RLOO (Ahmadian et al., 2024). The reward function is rule-based (DeepSeek-AI, 2025). For PPO, the critic model is initialized from Qwen2.5-0.6B. The learning rate for critic training is $1\times 10^{-5}$, the k3 KL-regularization coefficient is 0.001, and the GAE parameters $\lambda$ and $\gamma$ are both set to $1$. For DAPO, no KL regularization is applied; the high and low clip ratios are $0.28$ and $0.20$, respectively, and $L_{cache}$ is $1,024$. Across all algorithms, each training sample is sampled eight times ($K=8$). The actor is optimized with AdamW (Loshchilov & Hutter, 2017) at a learning rate of 1e-6, the training batch size is $128$, and training proceeds in a purely on-policy manner for up to 50 steps. All experiments are run on 8 $\times$ 96 GB NVIDIA H20 GPUs.
Appendix D More Experiments